OpenStudy (anonymous):

Will give medal Which equation demonstrates the associative property? A. –(33 + 4) = –33 + (–4) B. (–16 + 9) + (–19) = –16 + (9 + (–19)) C. (–5 + 11) + 3 = (11 + (–5)) + 3 D. 2 + (–24) = –24 + 2

OpenStudy (chosenmatt):

oh hold on did this type of questions before

TheSmartOne (thesmartone):

www.purplemath.com/modules/numbprop.htm

OpenStudy (anonymous):

i think it is C but not sure

OpenStudy (anonymous):

its b

OpenStudy (chosenmatt):

it is c

OpenStudy (anonymous):

wait b or c?

OpenStudy (anonymous):

(2+3)+4=2+(3+4)=9 \,

OpenStudy (anonymous):

b

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

It would be B not C..

OpenStudy (anonymous):

okay thanks

OpenStudy (secret-ninja):

Associative Property means- Of or relating to the property of an operation, such as addition or multiplication, which states that the grouping of numbers undergoing the operation does not change the result. AKA- You can add or multiply regardless of how the numbers are grouped. By 'grouped' we mean 'how you use parenthesis'. In other words, if you are adding or multiplying it does not matter where you put the parenthesis.

OpenStudy (anonymous):

yup right smart one

OpenStudy (secret-ninja):

Knowing this, what do you think the answer is?

OpenStudy (anonymous):

b

OpenStudy (chosenmatt):

thang!it gotta be c..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i thin im gonna give the medal to beo because he helped me corret the answer

TheSmartOne (thesmartone):

Who is beo

TheSmartOne (thesmartone):

Oh nevermind..

OpenStudy (anonymous):

@beomsu.c.59131

OpenStudy (anonymous):

him

OpenStudy (anonymous):

keep it up lol

TheSmartOne (thesmartone):

I know :)

OpenStudy (anonymous):

thanks yall

OpenStudy (secret-ninja):

No problem. :) Good job guys.

OpenStudy (anonymous):

it's c

TheSmartOne (thesmartone):

kind of late ^^^

OpenStudy (anonymous):

i guess